# Scaling Laws for Uncertainty in Deep Learning

### v1

@@ Abstract
Deep learning has recently revealed the existence of scaling laws, demonstrating that model performance follows predictable trends based on dataset and model sizes. Inspired by these findings and fascinating phenomena emerging in the over-parameterized regime, we examine a parallel direction: do similar scaling laws govern predictive uncertainties in deep learning? In identifiable parametric models, such scaling laws can be derived in a straightforward manner by treating model parameters in a Bayesian way. In this case, for example, we obtain $O(1/N)$ contraction rates for epistemic uncertainty with respect to the number of data $N$. However, in over-parameterized models, these guarantees do not hold, leading to largely unexplored behaviors.
In this work, we empirically show the existence of scaling laws associated with various measures of predictive uncertainty with respect to dataset and model sizes. Through experiments on vision and language tasks, we observe such scaling laws for in- and out-of-distribution predictive uncertainty estimated through popular approximate Bayesian inference and ensemble methods. Besides the elegance of scaling laws and the practical utility of extrapolating uncertainties to larger data or models, this work provides strong evidence to dispel recurring skepticism against Bayesian approaches: “In many applications of deep learning we have so much data available: what do we need Bayes for?” Our findings show that “so much data” is typically not enough to make epistemic uncertainty negligible.
@@ 1 Introduction
Figure 1: Deep learning uncertainty is predictable, empirically. ResNet-18 Epistemic Uncertainty scaling with the number of training data $N$ on CIFAR-10.
@@
Deep learning has recently revealed empirical scaling laws: test performance tends to scale by a power-law $f(x)\propto x^{-\gamma}$ of the size of the function, data or compute (Kaplan et al., 2020; Hoffmann et al., 2022). A related ‘double descent’ effect shows that increasing model capacity can surprisingly improve generalization (Belkin et al., 2019). These findings suggest that bigger models often perform better. A parallel question is: do similar scaling laws govern uncertainty in deep learning?
Bayesian deep learning provides a principled framework for quantifying uncertainty in neural networks by marginalizing weights, rather than relying on point estimates, to obtain predictive uncertainties. (Neal, 1996; MacKay, 1995). Recent contributions show that, under rather simple conditions, inferring only a subset of model parameters can retain the capacity to represent any predictive distributions, at a significantly lower cost (Sharma et al., 2023)
@@
The ability to produce well-calibrated predictive distributions and uncertainties is essential in decision-making and safety-critical applications, such as medical diagnosis and autonomous driving, where sound quantification of uncertainty can guide downstream choices (Papamarkou et al., 2024). However, exact Bayesian inference in large networks is intractable, which has motivated scalable approximations such as those based on the Laplace method (Ritter et al., 2018; Antoran et al., 2022), variational inference (Graves, 2011), Monte Carlo (mc) dropout (Gal and Ghahramani, 2016), along with (non-Bayesian) Deep Ensembles (Lakshminarayanan et al., 2017), which have proven effective and practical in obtaining sensible predictive uncertainties. Sampling-based methods, such as Stochastic Gradient Langevin Dynamics (sgld) (Welling and Teh, 2011) and Stochastic Gradient Hamiltonian Monte Carlo (sghmc) (Chen et al., 2014) provide scalable Markov chain Monte Carlo (mcmc) techniques for posterior inference over neural network weights. In the remainder of this paper, we will generally refer to Uncertainty Quantification (uq) methods to indicate Bayesian and non-Bayesian approaches designed to obtain predictive uncertainties.
@@
Despite recent criticism (Wimmer et al., 2023), current trends have focused on information-theoretic decomposition of predictive uncertainties into aleatoric and epistemic parts (Hüllermeier and Waegeman, 2021). In this work, we study how predictive uncertainties evolves in deep models as training data size $N$ grows. We evaluate a range of uq methods and architectures for both vision and language modalities. We systematically vary optimization settings, prior choices, stochasticity level, and inference techniques. We demonstrate that different configurations exhibit a range of scaling behaviors across uncertainty metrics, but crucially all of them follow power-laws.
While classic results show $O(1/N)$ posterior-variance contraction in identifiable parametric models (Cam, 1953), over-parameterized neural networks present challenges due to singular likelihoods, non-identifiability, and non-Gaussian posteriors. Although empirical scaling laws apply to test performance (Kaplan et al., 2020), a systematic understanding of how uncertainty scales with $N$, model size $P$, or optimization choices is lacking. Both Bayesian and non-Bayesian uq approaches yield ensembles, where Epistemic Uncertainty (eu) quantifies ensemble diversity in predictions. We are particularly interested in determining when eu vanishes leading to ensemble collapse; this is where we can leverage the practical utility of scaling laws.
Several theories characterize neural networks in the large-data limit. Singular Learning Theory (slt) models deep networks as singular statistical manifolds, claiming that singularities of loss minima determine learning behavior, and that generalization follows a power-law $\mathcal{L}-\mathcal{L}_{0}=\frac{\lambda}{n^{\gamma}}$ (Watanabe, 2009). Other perspectives include the manifold dimensionality hypothesis, linking generalization to the intrinsic dimensionality of learned representations (Ansuini et al., 2019), and mechanistic interpretability efforts explaining simple scaling phenomena via emergent internal representations (Nanda et al., 2023). We derive a formal connection of generalization and total uncertainty in the case of linear models. We believe that such a framework may help to explain the scaling behavior of uncertainty in deep models.
@@
(ii) Scaling patterns. We empirically demonstrate that predictive uncertainties evaluated on in- and out-of-distribution, follow power-law trends with the dataset size. This allows us to extrapolate to large dataset sizes and to identify data regimes where uq approaches remain relevant to characterize the diversity of the ensemble to a given numerical precision.
(iii) Theoretical insights. We derive a formal connection between generalization error in slt and total uncertainty in linear models. This novel analysis provides an interesting lead to explain the scaling laws observed in the experiments for over-parameterized models.
@@ 2.1 Uncertainty Quantification
Predictive uncertainty refers to metrics associated with an ensemble of predictive distributions and it can be decomposed into Aleatoric Uncertainty (au), arising from intrinsic data variability, and eu, reflecting uncertainty due to limited data or model knowledge (Hüllermeier and Waegeman, 2021).
The Total Uncertainty (tu) is the entropy of the mean predictive distribution
@@
In this work we study the power-law of test predictive uncertainty against training size $N$,
$\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\mathrm{Unc}({{\boldsymbol{\mathbf{x}}}}_{n}),\qquad\mathrm{Unc}\in\{\mathrm{TU},\mathrm{AU},\mathrm{EU}\}.$ (4)
While popular, these entropy-based metrics have limitations: (i) The standard decomposition of tu assumes additive separation of au and eu, which Wimmer et al. (2023) argues not to hold in complex deep models, and the difficulty of disentangling them hinders their interpretability (de Jong et al., 2024); (ii) Recent works highlight epistemic uncertainty collapse of large ensembles, leading to overly confident predictions (Kirsch, 2024; Fellaji and Pennerath, 2024).
@@ 2.2 Scaling Laws and Generalization
Empirical studies in deep learning have demonstrated that performance metrics scale predictably in model size, dataset size and compute. Initial findings by Hestness et al. (2017) and Kaplan et al. (2020) showed that test loss typically decreases following a power-law of the form
$\mathcal{L}(x)=\mathcal{L}_{\infty}+\left(\frac{x_{0}}{x}\right)^{\alpha}\,,$ (5)
where $x$ is the resource under analysis (e.g., dataset size $N$, model size $P$, or compute budget $C$), $\mathcal{L}_{\infty}$ is the irreducible loss, and $\alpha$ is modality- and task-specific ($x_{0}$ is a reference constant). These relationships hold consistently across model architectures, tasks, and modalities, as empirically shown by Henighan et al. (2020), who demonstrated similar loss scaling for language, image, video and multimodal domains. The scaling exponents differ across domains, but the functional form of the scaling law remains stable, indicating a general underlying behavior.
Theoretical explanations of scaling behaviors are only recently emerging: for example Sharma and Kaplan (2022); Bahri et al. (2024) linked $\alpha$ to the intrinsic dimension and spectral properties of the data, showing how the data geometry drives the reducible portion of the loss. However, despite these empirical and theoretical advances, a comprehensive understanding of how epistemic and aleatoric uncertainties scales with model and data size remain an open question. Indeed, a Bayesian perspective on uncertainty scaling laws is currently missing: it is unclear what kind of power laws uncertainty exhibits, if any. Understanding uncertainty scaling could significantly enhance the development and comprehension of Bayesian deep learning.
@@ 3 Methods
mc Dropout.
MC dropout is a simple and common inference method, with connections to variational inference (Gal and Ghahramani, 2016). During training standard dropout is applied, while at test time dropout masks are resampled to produce stochastic forward passes, yielding ensemble predictions. Due to its simplicity and universality, MC dropout is a good uq baseline.
@@
Gaussian are a common family of posterior approximations, and include Laplace approximations (Ritter et al., 2018) and Variational Inference (Graves, 2011). In Laplace approximations the curvature of the log-posterior at a local optimum determines the covariance of the Gaussian approximation. In variational inference, the optimization seeks an approximate distribution close to the true posterior in the Kullback-Leibler divergence (kl) sense.
@@
mcmc is the classic method of obtaining samples from the posterior over model parameters (Neal, 1996; MacKay, 1995). Gradient-based mcmc, such as Hamiltonian Monte Carlo, are some of the most effective samplers. Mini-batch-based sghmc (Chen et al., 2014) and Langevin dynamics (Welling and Teh, 2011) have been successfully proposed to sample from the posterior over parameters of deep neural networks of moderate size (Tran et al., 2022; Izmailov et al., 2021). We consider parameter sampling with both global parameter priors and Gaussian process functional priors (Tran et al., 2022).
@@
Deep ensembles (Lakshminarayanan et al., 2017) are a popular technique to obtain predictive uncertainties, despite lacking a full Bayesian interpretation. Multiple networks are trained independently from different seeds and uncertainty is estimated from the ensemble of predictions. Deep ensembles tend to provide better uncertainty estimates and are more robust to model misspecification than MC dropout. We train ensembles of size $M\in\{5,10\}$.
@@
We consider partially stochastic networks where only few layers are inferred, while rest are optimized in a standard way (Sharma et al., 2023).
@@ 4 Experiments
In this section, we present a detailed analysis of the extensive experimental results from our study. We begin by focusing on the most significant findings in the vision domain, highlighting the scaling behaviors of uncertainty both in- and out-of-distribution. Additionally, we report on an additional experiment conducted in the text classification setting with a GPT-2 language model. Our investigation spans a wide matrix of experimental configurations, aiming to explore a broad range of combinations across architectures, datasets and uq setups.
@@ 4.1 Image classification
We use common image classification architectures of ResNet (He et al., 2016), WideResNet (Zagoruyko and Komodakis, 2016) and Vision Transformer (ViT) (Dosovitskiy et al., 2021). We report CIFAR-10 and ImageNet32 results in the main paper, and results on CIFAR-100 in Appendix B.
4.1.1 ResNet and WideResNet architectures
@@
If not specified, we train these models for $400$ epochs with SGD optimizer (momentum $0.9$ and weight decay $5\times 10^{-4}$); in some experiments we adopt a cosine annealing scheduler on the learning rate and in others we do not. See Appendix C for more experimental details.
@@
Figure 2: ResNet and WideResNet($\mathbf{w}$-$\mathbf{d}$) uncertainty scaling on CIFAR-10 dataset: In (a) we use mc dropout with fixed dropout rate $p=0.2$ and $p=0.5$. In (b) we use Deep Ensembles (with $M=5$ and $M=10$ ensemble members). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training data. In (a), each point $\times$ corresponds to the average over $10$ independent folds (varying both data subsampling and model initialization), whereas in (b), the results for WideResNets are reported from a single fold. Dashed lines represent linear regressions fitted to the mean uncertainty metrics (au, eu, tu) on a fixed test set (see Section 2.1), following a power-law decay of the form $1/N^{\gamma_{TU}}$, $1/N^{\gamma_{AU}}$, and $1/N^{\gamma_{EU}}$. Both axes are shown on a logarithmic scale.
In ResNets, we add dropout layers after the convolutional blocks and before the fully connected layers following Kim et al. (2023). For WideResNets, we adopt the official implementation by Zagoruyko and Komodakis (2016), where dropout is applied between the two convolutional layers within each residual unit. We experiment with various dropout rates finding more expressivity in the scaling laws obtained with $p=0.5$, especially in ResNets architectures (Gal et al., 2017). Some results are reported in Fig. 2 and Fig. 7. Furthermore, we examine the impact of combining mc dropout with Sharpness Aware Minimization (sam) (Foret et al., 2020) to investigate how two techniques that independently improve generalization interact in terms of uncertainty. Our intuition behind the increasing trend of eu in Fig. 3 is that sam seeks to minimize a perturbed loss objective that penalizes sharp minima in the loss landscape. As dataset size increases, the curvature of the loss function (as measured by the spetral properties of the Hessian) increases. sam, which aims to avoid sharp valleys in the loss landscape, is therefore forced to navigate an increasingly curved landscape as $N\to\infty$. If the optimizer selects flatter basins in response to this increased curvature, it may avoid parameter regions that would otherwise exhibit low epistemic variance, and instead, it may prefer broader, flatter regions that span a more diverse set of functions—yielding higher epistemic uncertainty.
Figure 3: Impact of SGD+SAM on uncertainty scaling: ResNets on CIFAR-10 dataset - mc dropout ($p=0.5$).
@@
Across our experiments, we observe that the eu is typically smaller than the au. Finally, we report results training on CIFAR-10 and testing on CIFAR-10-C (corrupted), which allows us to explore how uncertainty behaves out-of-distribution (ood); results in Fig. 6.
@@
In Fig. 2, we report scaling laws for Deep Ensembles (Lakshminarayanan et al., 2017), using $M=5$ and $M=10$ independently trained models. Even for this uq approach, we observe the emergence of power-law scalings of uncertainties with respect to the number of data.
@@
We also observe a similar behavior for mcmc. In these experiments, we choose a weakly informative prior over all parameters with zero mean and standard deviation of $10$. In Appendix B, we also report results for mcmc after optimizing the priors according to the approach in Tran et al. (2022), where we target a Gaussian process (gp) with isotropic covariance with log-length-scale of $\frac{1}{2}\log{D}$ and and log-marginal variance of $2$. In Fig. 8 mcmc (1) and mcmc (2) refer to the approach of treating the first and first+second layers of the model in a stochastic fashion, while the other parameters are kept fixed to a pretrained solution.

### v2

@@ Abstract
Scaling laws in deep learning describe the predictable relationship between a model’s performance, usually measured by test loss, and key design choices such as dataset and model size. Inspired by these findings, we investigate a parallel direction: do similar scaling laws govern predictive uncertainties in deep learning? Identifiable models under a Bayesian framework exhibit $O(1/N)$ epistemic uncertainty contraction rate, with respect to dataset size $N$. However, these guarantees fail in over-parameterized regimes, where scaling behaviors remain largely unexplored. Across vision and language tasks, we show that in- and out-of-distribution predictive uncertainties, estimated via popular Bayesian and ensemble methods, follow predictable scaling laws relative to model and dataset size. Besides the elegance of scaling laws and the practical utility of extrapolating uncertainties to larger data or models, this work provides strong evidence to dispel recurring skepticism against Bayesian approaches: “In many applications of deep learning we have so much data available: what do we need Bayes for?” Our findings show that “so much data” is typically not enough to make epistemic uncertainty negligible.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Deep learning uncertainty is predictable, empirically. ResNet-18 Epistemic Uncertainty scaling with the number of training data NN on CIFAR-10.
Empirical scaling laws describe how deep learning performance improves predictably with increasing data, parameters, or compute, typically following a power-law decay
$f(x)\propto x^{\gamma},\quad\gamma<0$ (1)
@@
where $x\in\{N,P,C\}$ denotes dataset size, parameter count, or compute budge (Kaplan et al., 2020; Hoffmann et al., 2022). A related ‘double descent‘ effect shows that increasing model capacity can surprisingly improve generalization (Belkin et al., 2019). These findings suggest that bigger models often perform better. A parallel question is: do similar scaling laws govern uncertainty in deep learning?
Bayesian Deep Learning
@@
Bayesian deep learning provides a principled framework for quantifying uncertainty in neural networks by marginalizing weights, rather than relying on point estimates, to obtain predictive uncertainties (Neal, 1996; MacKay, 1995). Recent contributions show that, under rather simple conditions, inferring only a subset of model parameters can retain the capacity to represent any predictive distribution, at a significantly lower cost (Sharma et al., 2023). The ability to produce well‐calibrated predictive distributions and uncertainties is essential in decision‐making and safety‐critical applications, such as medical diagnosis and autonomous driving, where sound quantification of uncertainty can guide downstream choices (Papamarkou et al., 2024). However, exact Bayesian inference in large networks is intractable, which has motivated scalable approximations such as those based on the Laplace method (Ritter et al., 2018; Antoran et al., 2022), Variational Inference (vi) (Graves, 2011; Shen et al., 2024), Monte Carlo Dropout (mcd) (Gal and Ghahramani, 2016), along with (non-Bayesian) Deep Ensembles (de) (Lakshminarayanan et al., 2017), which have proven effective and practical in obtaining sensible predictive uncertainties. Sampling-based methods, such as Stochastic Gradient Langevin Dynamics (sgld) (Welling and Teh, 2011) and Stochastic Gradient Hamiltonian Monte Carlo (sghmc) (Chen et al., 2014) provide scalable Markov chain Monte Carlo (mcmc) techniques for posterior inference over neural network weights. In the remainder of this paper, we will generally refer to Uncertainty Quantification (uq) methods to indicate Bayesian and non-Bayesian approaches designed to obtain predictive uncertainties.
Predictive Uncertainty and Scaling
@@
Despite recent criticism (Wimmer et al., 2023), current trends have focused on information-theoretic decomposition of predictive uncertainties into aleatoric and epistemic parts (Hüllermeier and Waegeman, 2021). In this work, we study how predictive uncertainties evolve in deep models as dataset size $N$ and model size $P$ grow, evaluating a range of uq methods and architectures for both vision and language modalities. With stronger emphasis on scaling with respect to $N$, we systematically vary optimization settings, prior choices, stochasticity level, and inference techniques, demonstrating that different configurations exhibit a range of scaling behaviors across uncertainty metrics, but crucially all of them follow power-laws (see, e.g., Figure˜1).
While classic results show $O(1/N)$ posterior-variance contraction in identifiable parametric models (Cam, 1953), over-parameterized neural networks present challenges due to singular likelihoods, non-identifiability, and non-Gaussian posteriors. We specifically study the Linearized Laplace Approximation (lla) in Appendix˜B and confirm that, under network linearization, the posterior variance scaling asymptotically approaches the $1/N$ trend.
Although empirical scaling laws apply to test performance (Kaplan et al., 2020), a systematic understanding of how uncertainty scales with $N$, model size $P$, or computational budget is lacking. Both Bayesian and non-Bayesian uq methods yield ensembles, with Epistemic Uncertainty (eu) reflecting prediction diversity. One of the interesting aspects of scaling laws for uncertainty is to identify when eu vanishes, indicating ensemble collapse.
Singular Learning Theory (slt)
Several theories characterize neural networks in the large-data limit. slt models deep networks as singular statistical manifolds and shows that singularities of loss minima determine learning behavior. In particular, for Bayesian learning it predicts that the expected generalization error $G_{n}$ (the expected excess loss when going from $n$ to n+1n{+}1 training samples) decays as $\mathbb{E}[G_{n}]=\frac{\lambda}{n}+o\!\left(\frac{1}{n}\right)$, where $\lambda$ is the Real Log Canonical Threshold (rlct) (learning coefficient) (Watanabe, 2009). This last quantity governs the rate at which the generalization loss approaches the irreducible loss as $n$ (number of training data points) grows. We derive a formal connection between the generalization error and the Total Uncertainty (tu) in the case of linear models. We believe that such a framework may help explaining the scaling behavior of uncertainty in deep models.
Other perspectives include the manifold dimensionality hypothesis, linking generalization to the intrinsic dimensionality of learned representations (Ansuini et al., 2019), and mechanistic interpretability efforts explaining simple scaling phenomena via emergent internal circuits (Nanda et al., 2023).
Contributions
@@
(ii) Scaling patterns. We empirically demonstrate that predictive uncertainties evaluated on in- and out-of-distribution, follow power-law trends with dataset and model size. This allows us to extrapolate to large dataset sizes and to identify data regimes where uq approaches remain relevant to characterize the diversity of the ensemble to a given numerical precision.
(iii) Theoretical insights. We derive a formal connection between generalization error in slt and tu in linear models. This novel analysis provides an interesting lead to explain the scaling laws observed in the experiments for over-parameterized models.
@@ 2.1 Uncertainty Quantification
Predictive uncertainty refers to metrics associated with an ensemble of predictive distributions and it can be decomposed into Aleatoric Uncertainty (au), arising from intrinsic data variability, and Epistemic Uncertainty (eu), reflecting uncertainty due to limited data or model knowledge (Hüllermeier and Waegeman, 2021).
The tu is the entropy of the mean predictive distribution:
@@
In this work, we study the power-law of test predictive uncertainty, averaged over all the (fixed) test samples:
$\frac{1}{N_{\mathrm{test}}}\sum_{n=1}^{N_{\mathrm{test}}}\mathrm{U}({{\boldsymbol{\mathbf{x}}}}_{n}),\qquad\mathrm{U}\in\{\mathrm{TU},\mathrm{AU},\mathrm{EU}\}.$ (5)
Limitations of au and eu
While popular, these entropy-based metrics have limitations: (i) The standard decomposition of tu assumes additive separation of au and eu, which Wimmer et al. (2023) argues not to hold in complex deep models, and the difficulty of disentangling them hinders their interpretability (de Jong et al., 2025); (ii) Recent works highlight eu collapse of large ensembles, leading to overly confident predictions (Kirsch, 2024; Fellaji and Pennerath, 2024).
@@ 2.2 Scaling Laws and Generalization
Empirical studies in deep learning have demonstrated that performance metrics scale predictably with model size, dataset size and compute. Initial findings by Hestness et al. (2017) and Kaplan et al. (2020) showed that test loss typically decreases following a power-law of the form
$\mathcal{L}(x)=\left(\frac{x_{0}}{x}\right)^{\alpha}+\mathrm{const.}\,$, (6)
where $x$ is the resource under analysis (e.g., dataset size $N$, model size $P$, or compute budget $C$) and $\alpha$ is modality- and task-specific ($x_{0}$ is a reference constant). These relationships hold consistently across model architectures, tasks, and modalities, as empirically shown by Henighan et al. (2020), who demonstrated similar loss scaling for language, image, video and multimodal domains. The scaling exponents differ across domains, but the functional form of the scaling law remains stable, indicating a general underlying behavior.
Theoretical explanations of scaling behaviors are only recently emerging: for example Sharma and Kaplan (2022); Bahri et al. (2024) linked $\alpha$ to the intrinsic dimension and spectral properties of the data, showing how the data geometry drives the reducible portion of the loss. However, despite these empirical and theoretical advances, a comprehensive understanding of how eu and au scale with model and data size remain an open question. Indeed, a Bayesian perspective on uncertainty scaling laws is currently missing: it is unclear what kind of power laws uncertainty exhibits, if any. Understanding uncertainty scaling could significantly enhance the development and comprehension of Bayesian deep learning.
@@ 3 Methods
Monte Carlo Dropout.
mcd is a simple and common inference method, with connections to vi (Gal and Ghahramani, 2016). During training, standard dropout is applied, while at test time dropout masks are resampled to produce stochastic forward passes, yielding ensemble predictions. Due to its simplicity and universality, mcd is a good uq baseline.
@@
Gaussian are a common family of posterior approximations, encompassing both Laplace methods (Ritter et al., 2018) and variational inference (vi) (Graves, 2011). While Laplace relies on the local curvature of the log-posterior to define the Gaussian covariance, vi instead directly optimizes an approximate distribution to match the true posterior in the Kullback-Leibler divergence (kl) sense. Building on this line of work, we also include experiments with the Improved Variational Online Newton (ivon) optimizer (Shen et al., 2024), which unifies natural-gradient vi (Khan and Rue, 2023) with an online Newton method to achieve efficient large-scale Bayesian learning.
@@
mcmc is the classic method of obtaining samples from the posterior over model parameters (Neal, 1996; MacKay, 1995). Gradient-based mcmc, such as Hamiltonian Monte Carlo (hmc), are some of the most effective samplers. Mini-batch-based sghmc (Chen et al., 2014) and Langevin dynamics (Welling and Teh, 2011) have been successfully proposed to sample from the posterior over parameters of deep neural networks of moderate size (Tran et al., 2022; Izmailov et al., 2021). We consider parameter sampling with both global parameter priors and Gaussian process functional priors (Tran et al., 2022).
@@
de (Lakshminarayanan et al., 2017) are a popular technique to obtain predictive uncertainties, despite lacking a fully Bayesian interpretation. Multiple networks are trained independently from different seeds and uncertainty is estimated from the ensemble of predictions. de tend to provide better uncertainty estimates and are more robust to model misspecification than mcd. We train ensembles of size $M\in\{5,10\}$.
@@
We consider partially stochastic networks where only few layers are inferred, while the rest are optimized in a standard way (Sharma et al., 2023).
@@ 4 Experiments
In this section, we provide a detailed analysis of our experimental results. We first highlight the most significant findings in the vision domain, focusing on the scaling behavior of uncertainty both in-distribution and out-of-distribution (ood) (Section˜4.1). This is complemented by an analysis of uncertainty extrapolation, detailed in Figure˜15. We report an experiment in the text classification setting using a GPT-2 language model (Section˜4.2). For completeness, we also include a toy experiment with a Bayesian multilayer perceptron (mlp) (Section˜4.3). Our study covers a wide range of configurations across architectures, datasets, and uq setups, providing a comprehensive view of scaling trends.
@@ 4.1 Image classification
We use common image classification architectures of ResNet (He et al., 2016), WideResNet (Zagoruyko and Komodakis, 2016) and Vision Transformer (ViT) (Dosovitskiy et al., 2021). We report the CIFAR-10 results and part of the ImageNet32 and CIFAR-100 results in the main paper, while additional analyses are provided in Appendix˜C.
4.1.1 ResNet and WideResNet
@@
If not specified, we train these models for $400$ epochs with Stochastic Gradient Descent (sgd) optimizer (momentum $0.9$ and weight decay $5\times 10^{-4}$); in some experiments we adopt a cosine annealing scheduler on the learning rate and in others we do not. See Appendix˜D for more experimental details.
Monte Carlo Dropout
@@
Figure 2: ResNet-$\mathbf{d}$ uncertainty scaling on CIFAR-10 and CIFAR-100 datasets: We use mcd (with fixed dropout rate $p=0.2$ and $p=0.5$); each point $\times$ corresponds to the average over $10$ independent folds (varying both data subsampling and model initialization). We consider $25\%$, $50\%$, $75\%$ and $100\%$ subsets of the training dataset. Dashed lines represent linear regressions fitted to the mean uncertainty metrics (au, eu, tu) on a fixed test set (see Section˜2.1), following a power-law decay of the form $N^{\gamma_{TU}}$, $N^{\gamma_{AU}}$, and $N^{\gamma_{EU}}$. Both axes are shown on a logarithmic scale.
In ResNets, we add dropout layers after the convolutional blocks and before the fully connected layers following Kim et al. (2023). For WideResNets, we adopt the official implementation by Zagoruyko and Komodakis (2016), where dropout is applied between the two convolutional layers within each residual unit. We experiment with various dropout rates finding more expressivity in the scaling laws obtained with $p=0.5$, especially in ResNets architectures (Gal et al., 2017). Some results are reported in Figure˜1, Figure˜2, and Figure˜6.
Across our experiments, we observe that the eu is typically smaller than the au. Regarding the decreasing trend of au that we observe across various experiments, Wimmer et al. (2023) also report that au can decrease under limited data, making estimates potentially unreliable in such regimes. Prior work (de Jong et al., 2025; Mucsányi et al., 2024) further shows that au and eu often correlate in image classification, suggesting entanglement in practice.
Monte Carlo Dropout combined with Sharpness Aware Minimization
Figure 3: Impact of sgd +sam on uncertainty scaling: ResNets on CIFAR-10 dataset using mcd ($p=0.5$) for uq. sam biases solutions towards flatter minima and the combination with mcd preserves functional diversity as data size increases.
We also study mcd combined with sam (Foret et al., 2020) to assess how two generalization methods interact in terms of uncertainty. We hypothesize that the increasing eu in Figure˜3 arises because sam, by avoiding sharp minima, is forced into flatter basins as $N\to\infty$: by selecting flatter basins under increased curvature, the optimizer may avoid regions of low epistemic variance and instead favor broader regions spanning more diverse functions. This aligns with recent findings that sam can enhance ensemble diversity by trading off sharpness minimization with data-driven de-correlation (Lu et al., 2024), while mcd further promotes functional diversity via implicit bagging.
Figure 4: ResNets uncertainties out-of-distribution: We use mcd with $p=0.2$ in (a) and $p=0.5$ in (b). For models trained on incrementally larger training subsets of CIFAR-10, we report the predictive uncertainties when testing on the (whole) CIFAR10-C dataset, averaged over all corruption levels (11-55) and corruption types considered. ood, we expect to observe larger uncertainties - eu, for instance, should decay gradually as the data space becomes increasingly populated with additional samples within the same domain (in-fill).
Uncertainty out-of-distribution
@@
We report results training on CIFAR-10 and testing on CIFAR-10-C (corrupted), which allows us to explore how uncertainty behaves ood; results in Figure˜4.
Figure 5: WideResNet -$\mathbf{w}$-$\mathbf{d}$ uncertainty scaling on CIFAR-10 dataset: We consider WideResNet-40-4 and WideResNet-28-10 and perform uq using de with $M=5$ and $M=10$ ensemble members.
Deep Ensembles
@@
In Figure˜5, we report scaling laws for de (Lakshminarayanan et al., 2017), using $M\in\{5,10\}$ independently trained models. Even for this uq approach, we observe the emergence of power-law scalings of uncertainties with respect to the number of data.
Figure 6: Uncertainty estimates of ResNets on ImageNet-32 dataset: The first two subplots report mcd ($p=0.5$) uncertainties for ResNet-18 and ResNet-34, trained with fixed learning rate $10^{-3}$ on $9$ increasing subsets of the training data. We also report uncertainties obtained through mcmc considering only the first layer stochastic (mcmc (1)) and only the first two layers stochastic (mcmc (2)) for ResNet-18 trained on $4$ increasing subsets of the training data.
Markov Chain Monte Carlo
@@
We also observe a similar behavior for mcmc. In these experiments, we choose a weakly informative prior over all parameters with zero mean and standard deviation of $10$. In Appendix˜C, we also report results for mcmc after optimizing the priors according to the approach in (Tran et al., 2022), where we target a Gaussian process (gp) with isotropic covariance with log-length-scale of $\frac{1}{2}\log{D}$ and and log-marginal variance of $2$. In Figure˜6, mcmc (1) and mcmc (2) refer to the approach of treating the first and first+second layers of the model in a stochastic fashion, while the other parameters are kept fixed to a pretrained solution (Sharma et al., 2023).
ResNets uncertainty scaling with model parameters
Figure 7: Uncertainty scaling with model parameters of ResNets on CIFAR-10: We evaluate a wide spectrum of ResNet architectures on CIFAR-10 dataset. The corresponding parameter counts are: ResNet-18 ($11.7M$), ResNet-50 ($25.6M$), ResNet-101 ($44.5M$), and ResNet-152 ($60.2M$). Left: We train each model on the entire dataset using the ivon optimizer for $200$ epochs, following the setup of Shen et al. (2024) (details in Appendix˜D). Right: We compare against mcd ($p=0.5$), training for $400$ epochs with sgd and fixed learning rate $10^{-3}$.
Figure 8: ViT-small uncertainty training dynamics on CIFAR-10 dataset: In (a) we use mcd (fixed $p=0.5$ both in the embeddings and in the transformer blocks) and we train the model for $200$ epochs with Adam optimizer (Kingma and Ba, 2017) and cosine annealing. In (b) we train the same model for $1000$ epochs and fixed learning rate $10^{-4}$. The training dynamics show that the the shape/speed of convergence of of tu, au, eu depends on the optimization trajectories underneath.
We investigate uncertainty scaling with model size in a preliminary experiment using mcd and the ivon optimizer. ivon shows the expected increase in eu with larger models, whereas mcd does not, likely reflecting limitations of the inference scheme. Parameter permutation symmetries (e.g., swapping hidden units or attention heads) generate exponentially many modes that yield the same function. Recent works (Rossi et al., 2023; Laurent et al., 2023; Gelberg et al., 2024) show these modes are connected by low-loss paths and are not functionally distinct. Thus, increasing model capacity adds redundancy rather than distinct hypotheses, explaining the weak or flat eu dependence observed in Figure˜7.